OpenStudy (anonymous):

Find the factors . x^2+2x-24

pooja195 (pooja195):

Look at the equation in slandered form \[\huge~\rm~\bf~ax^2+bx+c\] \[\huge~\rm~\bf~\color{purple}{x^2+2x-24}\] 

pooja195 (pooja195):

 What TWO numbers multiply to make -24 but add to make 2 ?

OpenStudy (kayders1997):

(x-4)(x+6) the reason why this works is because 6 times 4 is -24 and than when you add 6 and 4 it will equal to 2 because both the terms are getting multiplied by x and than x times x is x^2

OpenStudy (kayders1997):

what?

OpenStudy (kayders1997):

x would equal 4 and -6

OpenStudy (tkhunny):

Still trying to figure out what "slandered" form is. :-) Excellent typo. If you are factoring this: x^2+2x-24 It MUST look like this: (x + ____)(x - ____) The blanks MUST be factors of 24. 1*24 = 24 2*12 = 24 3*8 = 24 4*6 = 24 You mist find the pair with a difference of 2. What I'm trying to say is that you don't have to guess and you don't have to wander into bewilderment. Write down what you know. Think about what is not yet known.

pooja195 (pooja195):

My typos XD

OpenStudy (tkhunny):

"you mist find" -- no one is immune.